OpenStudy (anonymous):

Find the equation of the plane with the given description: Contains the lines r1(t)=<2,1,0>+ and r2(t)=<2,1,0>+<3t,t,8t> help? :( @ganeshie8 @perl

ganeshie8 (ganeshie8):

You need a `point` and `normal vector` to write teh equation of plane

OpenStudy (anonymous):

are you telling me the steps or? because thats all the problem gave me ._.

ganeshie8 (ganeshie8):

I'll give a hint : cross product of direction vectors of given lines

OpenStudy (anonymous):

ok. hold on...

OpenStudy (anonymous):

-_- is it 13x-1y-5z=d?

ganeshie8 (ganeshie8):

normal vector is <13, 1, -5> use any points on the plane and see jf you can write out the equation

OpenStudy (anonymous):

i checked the answer and it is however, it equaled 13x-y-5z=27 to find that 27 id have to find a random point? how?

ganeshie8 (ganeshie8):

you should get 13x\(\color{red}{+}\)y-5z=d

OpenStudy (anonymous):

oh yeah, +y sorry

ganeshie8 (ganeshie8):

you can find the value of "d" by plugging in any point (you're given two lines, think a bit.. )

ganeshie8 (ganeshie8):

aren't you given two ready made points as well ? r1(t)= `<2,1,0>`+<t,2t,3t> and r2(t)= `<2,1,0>`+<3t,t,8t>

OpenStudy (anonymous):

yeah ._.

ganeshie8 (ganeshie8):

you may use that point

ganeshie8 (ganeshie8):

13x+1y-5z=d plugin the point (2, 1, 0) and find out the value of "d"

OpenStudy (anonymous):

ohhh! gosh youre better at explaining and teaching than my UCLA professors.